Given are three vectors. Vector

A

of magnitude 5.30 units is in a direction θ = 36.0° above the positive x-axis. Vector

B

of magnitude 5.00 units is directed in the direction of the positive y-axis. Vector

C

of magnitude 5.80 units is in a direction ϕ = 37.0° clockwise from the negative y-axis.

Find the magnitude and direction of 2

AB + 3C

magnitude     units
direction     ° counterclockwise from the +x-axis
Answer #1

A B f 83600 B = d 0:32 -у A5.3.2013c1 + 5.30514362 A=4.299 + 3012 5:00 Co-580sin3_5-80C0337 C = -3.489 - 9.64 24-1+30 = (2x9.